OpenStudy (pinknabastak):

a car is traveling at 16m/s when the brakes are applied. The brakes decelerate the car at the rate of 5.0m/s^2. How long will it take the car to stop?

jimthompson5910 (jim_thompson5910):

Use the formula \[\Large V_{f} = V_{o} + a*t\] Vf is the final velocity Vo is the initial velocity a = acceleration t = time

jimthompson5910 (jim_thompson5910):

In this case Vf = 0 (we want the car to be stopped) Vo = 16 a = 5 t = unknown (you're solving for this)

OpenStudy (pinknabastak):

wouldnt it be -5??

OpenStudy (pinknabastak):

because it decelerating?

OpenStudy (pinknabastak):

not time the acceleration

jimthompson5910 (jim_thompson5910):

yes sorry, a = -5

OpenStudy (pinknabastak):

ok... so its 0=16+(-5)(t)

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (pinknabastak):

-16=-5t

OpenStudy (pinknabastak):

so 3.2=t

jimthompson5910 (jim_thompson5910):

getting the same

OpenStudy (pinknabastak):

is it 3.2 seconds

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (pinknabastak):

hahaaha

OpenStudy (pinknabastak):

OMG THANK YOU!!!! ok what about this......you want to put a new cd in your car stereo. while driving at 45 mi/h, you look away for 2.0s. What distance (in miles) do you "drive blind"?

OpenStudy (skullpatrol):

all yours @jim_thompson5910 :D

OpenStudy (pinknabastak):

haha :P @skullpatrol

jimthompson5910 (jim_thompson5910):

you would use the formula x = v*t x = distance traveled v = velocity t = time

jimthompson5910 (jim_thompson5910):

be sure to convert 2 seconds to hours or convert 45 mph to mi/sec

OpenStudy (pinknabastak):

uh......what? :/

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (pinknabastak):

converting that and howd you come up with that equaation?

jimthompson5910 (jim_thompson5910):

it's similar to the Distance = Rate * Time equation or D = r*t equation

jimthompson5910 (jim_thompson5910):

but with physics, they use x = v*t same idea, just different variables

jimthompson5910 (jim_thompson5910):

how many seconds are in a minute?

OpenStudy (pinknabastak):

ok so d=vt

OpenStudy (pinknabastak):

60

jimthompson5910 (jim_thompson5910):

yes that's another way to think of it

jimthompson5910 (jim_thompson5910):

and how many minutes in an hour?

OpenStudy (pinknabastak):

60

jimthompson5910 (jim_thompson5910):

so there are 60*60 = 3600 seconds in 1 hour

OpenStudy (pinknabastak):

ok yes

jimthompson5910 (jim_thompson5910):

1 hour = 3600 seconds how can you use that to convert 2 seconds to hours?

OpenStudy (pinknabastak):

do 2/3600

OpenStudy (pinknabastak):

0.0006

jimthompson5910 (jim_thompson5910):

correct, which is 1/1800 = 0.000555 roughly

jimthompson5910 (jim_thompson5910):

a very small piece of an hour

jimthompson5910 (jim_thompson5910):

now you can compute d = v*t

jimthompson5910 (jim_thompson5910):

v = 45 mph t = 0.000555 hours

OpenStudy (pinknabastak):

0.025

OpenStudy (pinknabastak):

miles

OpenStudy (skullpatrol):

Great work!

jimthompson5910 (jim_thompson5910):

0.024975 ---> 0.025 looks good

OpenStudy (pinknabastak):

yay!!! thanks a bunch!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (pinknabastak):

i have one more question if thats ok with you?

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (pinknabastak):

ok so A typical world-class male sprinter will accelerate to his maximum velocity in 4.0s. If the runner covers a distance of 55m in that time, what is his acceleration?

OpenStudy (pinknabastak):

d=55 t=4 a=? but what equation should i use?

jimthompson5910 (jim_thompson5910):

here's a good chart to have as reference http://www.studyphysics.ca/newnotes/20/unit01_kinematicsdynamics/chp04_acceleration/images/formulas.GIF

jimthompson5910 (jim_thompson5910):

one sec, the formula isn't on that page, but I'll try to find it

OpenStudy (pinknabastak):

haha yea...it would be but i dont know the velocity...

jimthompson5910 (jim_thompson5910):

I wonder if the initial velocity is 0?

jimthompson5910 (jim_thompson5910):

if so, then we can use the second formula on that link I posted

jimthompson5910 (jim_thompson5910):

If initial velocity is 0, then d = 55 Vi = 0 t = 4 use the second formula to solve for 'a'

OpenStudy (pinknabastak):

but we want to know the acceleration..

OpenStudy (pinknabastak):

ohhh ok

jimthompson5910 (jim_thompson5910):

\[\Large d = V_{i}*t + 1/2*at^2\] \[\Large 55 = 0*4 + 1/2*a*(4)^2\] solve for 'a'

OpenStudy (pinknabastak):

ok give me just a sec

OpenStudy (pinknabastak):

ok i got 1.72m/s^2

jimthompson5910 (jim_thompson5910):

that's too small

OpenStudy (pinknabastak):

uhhh......ok i did 55=0(4)+1/2(a)(4)^2 then 55=1/2(a)(16) then 27.2=a(16) then 1.72=a

jimthompson5910 (jim_thompson5910):

you made an error going from 55=1/2(a)(16) to 27.2=a(16)

jimthompson5910 (jim_thompson5910):

you should multiply both sides by 2, not divide

OpenStudy (pinknabastak):

....wow im retarded

jimthompson5910 (jim_thompson5910):

nah you're not, silly mistake really

OpenStudy (pinknabastak):

ok lol gimme a sec...

OpenStudy (pinknabastak):

ok i got 6.9 m/s^2

OpenStudy (pinknabastak):

that just shows that it stoo late to be doing physics haha

jimthompson5910 (jim_thompson5910):

I'd go with 6.875

OpenStudy (pinknabastak):

ok thats no problem with me! :)

OpenStudy (pinknabastak):

Thanks a whole bunch! You really helped! I will definitely use that formula table throughout this year! Thanks againn!

jimthompson5910 (jim_thompson5910):

you're welcome